OpenStudy (anonymous):

Graph the inequality. 4x – 2y ≤ -12

OpenStudy (anonymous):

solve for y to get it in y=mx + b form

OpenStudy (anonymous):

gotcha so far

OpenStudy (anonymous):

so then

OpenStudy (anonymous):

graph it normally and if y => shade above if y <= shade below :)

OpenStudy (anonymous):

6+2*x <= y

OpenStudy (anonymous):

so just graph it like you would a y=mx+b equation and shade above :)

OpenStudy (precal):